You buy a share of The GLS Corporation stock for the market price, P0.  You expect it to pay dividends each year, including a dividend of D3 in Year 3, and each of the dividends has grown at the historical constant growth rate.  You expect to sell it at a price P3 at the end of three years.  Calculate the growth rate, the expected dividend yield, and the stock's expected annual total rate of return on the stock. In other words, the annualized total return you would expect to earn on the stock each year (total yield). The Total Yield is also known as the Total Return.
INPUT DATA D3 $1.46 P0 $23.25 P3 $28.55
Growth Rate
Dividend yield
Total yield
Answer #1

1)Growth Rate =[P3/P0]^1/n - 1

         =[28.55/23.25]^1/3    -1

        = [1.22796]^.33333     -1

       = 1.0708-1

       = .0708 or 7.08%

2)D3 =D0(1+g)^n

   1.46 = D0(1+.0708)^3

    1.46 = D0 * 1.22779

     D0 = 1.46/1.22779

           = 1.18913

Dividend yield = D0(1+G) /P0.

                       = 1.18913(1+.0708)/ 23.25

                      = 1.27332/23.25

                      = .0548 or 5.48%

Total yield = Dividend yield +growth

             = 5.48+7.08

            = 12.56%